Split funnels were given to Oemer for cleaning. Have received them back after lunch.
Valve support is installed again (see picture 1).
Ion pump was connected but not activated. Michael will help in doing so. No vacuum support to help will RGA. Will proceed without it.
Neg and ion cables were labeled (see pictures 2 and 3).
Chamber was vented using nitrogen.
|